OpenStudy (greatlife44):

Calculus 1 question: Find the area of the largest rectangle that can be inscribed in a circle of radius 1.

OpenStudy (greatlife44):

\[x^{2} + y^{2} = 1\]

OpenStudy (ace-n-it):

k

OpenStudy (greatlife44):

Here's the circle 

OpenStudy (greatlife44):

then I know that a rectangle we can express the area as l*w which in terms of the circle can be xy = A

OpenStudy (greatlife44):

@ganeshie8 @pooja195

OpenStudy (greatlife44):

@mayankdevnani

OpenStudy (greatlife44):

You see @surjihayer I also thought we could express the length like this: \[\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}} = \sqrt{4} = 2x \] \[[\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}} = \sqrt{4} = 2y \] A = 2x*2y

OpenStudy (fortytherapper):

I'm not completely sure, but I may have an idea Since we are fitting a rectangle in a circle, then that if we draw a diagonal from the center to a corner, that would be the hypotenuse and that would equal the radius, right? So if we created a triangle with that hypotenuse, maybe we can find the length some way? What would the length also be equal to?

OpenStudy (fortytherapper):

Actually, I should read replies before I actually start answering. I think you took the same path

OpenStudy (greatlife44):

@fortytherapper @zelda here is what I was thinking \[x^{2}+y^{2} = 1 \] \[y^{2} = 1-x^{2}\] \[y = \sqrt{1-x^{2}}\] 4xy = Area \[4x \sqrt{1-x^{2}} = Area \]

OpenStudy (greatlife44):

then I think we need to get the derivative

OpenStudy (fortytherapper):

Would this derivative involve 2 product rules and a chain rule?

OpenStudy (greatlife44):

Yeah we need to use a product and a chain rule... \[\frac{ dA }{ dx }(4x \sqrt{1-x^{2}}) \] \[4x \sqrt{1-x^{2}}' + 4x' \sqrt{1-x^{2}} \]

OpenStudy (greatlife44):

\[4x(0.5)*(1-x^{2})^{-1/2} * -x + 4 \sqrt{1-x^{2}}\] Then we get this \[\frac{ -4x^{2} }{ \sqrt{1-x^{2}} } +4\sqrt{1-x^{2}}\] \[-4x^{2}+4(1-x^{2}) = -4x^{2}+4-4x^{2} = -8x^{2} = -4\] \[\sqrt{ x^{2} }= \sqrt{\frac{ -4 }{ -8 }} = \frac{ 1 }{ \sqrt{2} }\] then we would plug this back into our original equation \[4(\frac{ 1 }{ \sqrt{2} })*\sqrt{1-(\frac{ 1 }{ \sqrt{2} })^{2}} = 2 \]

OpenStudy (greatlife44):

I asked my friend before, I didn't know how to set it up the right equation to find the area. but the rest isn't bad

OpenStudy (fortytherapper):

You're a genius, nice!